levy:

Write the equation of the circle with center (−3, 2) and (6, 4) a point on the circle.

dude:

So the general equation is \(x^2+y^2=r^2\) You want to add/subtract to the x to move it horizontally and add/subtract to the y value to move it vertically You want to move the x and y so that it is focused on the center (In this case, 3 to the left of y and 2 up) Moving is counterintuitive so you add 3 and subtract 2 \((x+3)^2+(y-2)^2=r^2\) To find the r value, find the distance between both points and solve Once you have the radius (which should be ~9.2) You should get 85 So your equation should be \((x+3)^2+(y-2)^2=85\) Answered solely because you were offline.